# The Brownian loop measure on Riemann surfaces and applications to length spectra

### v1

@@ 4.3 Zeta-regularized determinant of Laplacian
Acknowledgments. We would like to thank Jayadev Athreya, Frédéric Naud, Bram Petri, Steffen Rohde, Fredrik Viklund, Wendelin Werner, Yunhui Wu, and Anton Zorich for the helpful discussions and/or comments on an earlier version of the manuscript, and Minjae Park for pointing us to the reference [28]. This project is funded by the European Union (ERC, RaConTeich, 101116694) [^1] Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. .

### v2

@@ Appendix A Basics on Brownian motion on Riemannian manifolds
Acknowledgments. We would like to thank the anonymous referees for constructive comments, Jayadev Athreya, Frédéric Naud, Minjae Park, Bram Petri, Steffen Rohde, Fredrik Viklund, Wendelin Werner, Yunhui Wu, and Anton Zorich for the helpful discussions and/or comments on an earlier version of the manuscript, and Yves Le Jan for pointing us to the reference [31]. This project is funded by the European Union (ERC, RaConTeich, 101116694) [^2] Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. .